OpenStudy (anonymous):

Write the point slope form of the equation of the line described Through (4,2), parallel to y=-3/4x -5 Through (-3, -3) parallel to y=7/3x+3 Answer: y+3=7/3(x+3) Through (-4,0), parallel to y=3/4x-2 Answer: y-0=3/4(x+4) Are these correct :? @hartnn or anyone else Through (-1,4) parallel to y=-5x+2 Answer: y-4=-5(x-4)

OpenStudy (anonymous):

The first one y-2=-3/4(x-4) forgot to add the answer

OpenStudy (alexandervonhumboldt2):

checking your work..............

Miracrown (miracrown):

They are all correct except a small mistake in the last one

OpenStudy (alexandervonhumboldt2):

Through (-3, -3) parallel to y=7/3x+3 Answer: y+3=7/3(x+3) correct Through (-4,0), parallel to y=3/4x-2 Answer: y-0=3/4(x+4) correct Through (-1,4) parallel to y=-5x+2 Answer: y-4=-5(x-4) incorrect: the x-4 part is incorrect:

Miracrown (miracrown):

The point is (-1, 4), right? Did you use both the -1 and the 4 in the answer you wrote?

OpenStudy (anonymous):

Oh I see what I did wrong oops

Miracrown (miracrown):

We all make mistakes here and there. So don't worry, as long as you realize your mistake and learn from it that is what truly matters in the end. :)